Circular No. 6603

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1737–31 F. E. Marshall, Goddard Space Flight Center; and D. M. Smith, Uni-

versity of California, Berkeley, report for the RXTE team: “Recent scans over the position of the new x-ray transient GRS 1737–31 (IAUC 6599) with the Proportional Counter Array on the Rossi X-ray Timing Explorer reveal that the outburst began between Feb. 17.0 and 20.4 UT and peaked ∼ 2 weeks later. The following fluxes in milliCrabs (2–10 keV) were detected: Feb. 17.0, < 3; 20.4, 10; 27.1, 14; 27.9, 25; Mar. 6.1, 23; 9.9, 25; 11.0, 20; 19.9, 17. A pointed observation with the PCA and HEXTE was carried out during Mar. 21.69–21.73.”

4U 1812–12 L. Burderi, N. Robba, Istituto di Fisica, Universitá di Palermo; M.

Guainazzi, D. Ricci, M. Capalbi, G. Celidonio, A. Coletta, R. Ricci, and M. Smith, BeppoSax, Rome; and A. R. King, G. A. Wynn, and U. Kolb, Astronomy Group, University of Leicester, report: “On Mar. 12.220949 UT, the Wide Field Camera No. 1 onboard the BeppoSAX satellite detected an x-ray burst in the energy range 1.5–26.0 keV. The burst was seen in the ratemeter as a single peak, showing a moderately fast rise (of the order of 10 s) and an exponential decay lasting for ∼ 70 s (the decay time is of the order of 20 s). The burst intensity was 1.2 Crab in the band 1.5– 10.0 keV and 0.6 Crab in the band 10.0–26.0 keV. The derived position is α = 18^{h}14^{m}41^{s}.8, δ = −11^{o}52′17′′ (equinox 2000.0) with an estimated error radius of ∼ 18′. The error box contains the source 4U 1812–12 (α = 18^{h}15^{m}12^{s}.0, δ = −12^{o}04′59′′, equinox 2000.0), which is classified as a low-mass x-ray-binary bursting source. No other x-ray binary sources are contained within 60′ radius. This implies that the identification with 4U 1812–12 is quite robust.”

1E 1740.7–2942 AND GRO J1744–28 Further to the item concerning GRS 1737–31 on IAUC 6599, R. Sun-

yaev reports for the GRANAT/SIGMA Team: “Other sources detected in the galactic-center field during preliminary analysis of the data are 1E 1740.7–2942 (fluxes 96 ± 16 and 94 ± 19 mCrab in the bands 35–75 and 75–150 keV, respectively) and GRO J1744–28 (fluxes 107 ± 16 and 10 ± 18 mCrab, respectively). GRANAT will continue observations of the galacticcenter region until the end of March.”

1997 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green